What is the % pH change if a student added 0.0025 mol of Ca(OH)2 to a 1.00 L buffer solution containing 0.100 M HC2H3Oand 0.200 M NaC2H3O2.


1
Expert's answer
2021-04-05T07:59:54-0400

pH = pKa - lg"\frac{C(acid)}{C(salt)}"

pKa(HC2H3O) = 4.75

pH1 = 4.75 - "lg\frac{0.1}{0.2}" = 5.05

C2(acid) = 0.1 - 0.0025 "\times" 2 = 0.095 M

C2(salt) = 0.2 + 0.0025 "\times" 2 = 0.205 M

pH2 = 4.75 - lg"\frac{0.095}{0.205}" = 5.86

"\Delta"pH = 5.86 - 5.05 = 0.81

%pHchange = "\frac{0.81 \times 100}{5.05}" = 16 %

Answer: 16 %.